# Debiasing surgeon: fantastic weights and how to find them

### v1

@@ Abstract
Nowadays an ever-growing concerning phenomenon, the emergence of algorithmic biases that can lead to unfair models, emerges. Several debiasing approaches have been proposed in the realm of deep learning, employing more or less sophisticated approaches to discourage these models from massively employing these biases. However, a question emerges: is this extra complexity really necessary? Is a vanilla-trained model already embodying some “unbiased sub-networks” that can be used in isolation and propose a solution without relying on the algorithmic biases? In this work, we show that such a sub-network typically exists, and can be extracted from a vanilla-trained model without requiring additional training. We further validate that such specific architecture is incapable of learning a specific bias, suggesting that there are possible architectural countermeasures to the problem of biases in deep neural networks.
Keywords Debiasing  $\cdot$ Pruning  $\cdot$ Freezed model  $\cdot$ Deep learning
@@ 1 Introduction
Multiple solutions are proposed in the last luster, all relying on relatively heavy re-training approaches. More specifically, in the realm of DL debiasing, we can identify three main lines of research: (i) supervised, where the labels of the bias are provided; (ii) bias-tailored, where proxy models can capture specific biases; (iii) unsupervised, where biases are guessed directed within the vanilla-trained model. All three approaches typically require heavy training procedures for either properly tuning the hyper-parameters of the models, or training multiple times the neural networks. This comes at another very relevant environmental cost, and besides it is not granted that such solutions can always apply to any architecture/task. Some recent studies hinder the possibility of reusing pre-trained models and removing bias sources with some adjustments [65, 66], but none of them provides a final solution yet.
In this work we show the existence of unbiased sub-networks in vanilla-trained models, providing also some analysis on the final performance. In a nutshell, we learn how to mask trained weights such that the information about the bias becomes non-extractable by the layer(s) entitled to solve the target task. Therefore, we Find the Fantastic Weights (FFW) constituting an unbiased sub-network that solves the tasks, without further fine-tuning (Fig. 1). Our results, besides providing a clear indication that such a sub-network exists in vanilla-trained overfitting models, contribute to the DL debiasing community, showing that these solutions are in principle achievable without relying on over-complex training schemes.
@@ 2 Related works
Spurious correlations. The risk of the insurgence of spurious correlations employed for a given target task is a broadly known and acknowledged issue in the DL community. This phenomenon is known as shortcut learning [23], where the DL algorithm relies on the simplest correlations to fit the training set, but do not generalize well. Some of these can range from the texture of images [24] to biases inherent in language [26], or even sensitive variables like ethnicity or gender [47, 21]. Such behavior raises practical concerns, particularly in applications where the reliability of deep networks is critical, such as healthcare, finance, and legal services [15].
@@
Besides these, a new class of unsupervised approaches is rising [43], where in principle little or even no clue on the nature of the bias is provided. Specifically, we can recognize models pre-capturing texture biases [59, 25], leveraging groups imbalances [39, 16, 3] and assuming that the bias is learned in early training stages [45, 37].
Studying impacts of neural architectures. Recently a lot of attention has been devoted in studying the influence of a deep neural network’s architecture on its generalizability. For example, Diffenderfer et al. [19] built a pruning algorithm inspired by the lottery ticket hypothesis [22] to craft compact and resilient network architectures. Similarly, Bai et al. [7] tackles a similar issue through the lenses of out-of-distribution performance. This class of approaches, although effective at the cost of enhanced computational complexity [54], does not entirely eradicate connections to spurious input attributes. On the other side of the coin, Zhang et al. and Zhao et al. suggest the efficacy of pruning weights associated with such attributes sided with some fine-tuning [64, 2], finding some counter-arguments in the literature [10] and even with works suggesting the need to employ corrective distillation terms [48]. The discussion around these empirical findings is still open.
In a recent work, it has been empirically showcased the presence of subnetworks within neural networks that exhibit reduced susceptibility to spurious features [64]. By capitalizing on the modular nature inherent in neural architectures [18], Zhang et al. were able to demonstrate that, in principle, it is possible to select, already at initialization, networks that will avoid the employment of biased features. However, this is not a sufficient condition to still guarantee the presence of such sub-networks in an already-trained model. In the era of foundation models, indeed, training a model from scratch can be very costly and impractical [63]. In this work we focus on this problem, characterizing the condition of extractability for such debiased sub-networks, drawing as well a link between bias removal and impact on the performance of the target task. We will sketch the theory motivating our approach, and we will study the problem of extracting these subnetworks, trained with vanilla strategies, without requiring finetuning.
@@ 3.1 Removing the bias impacts the performance
In this section, we will define the problem of learning a mapping between the input $\mathbf{x}$ and its target output $\hat{y}$ on a given dataset $\mathcal{D}$ (that is traditionally slit in train, validation, and test), under the assumption that there is an underlying bias $\hat{b}$ (namely, “spurious” correlations) associated to it. We can model this problem working with the random variables associated with these quantities: we define $\hat{Y}$ as the random variable associated with the target output, $Y$ as the random variable associated with the output of the classifier $\mathcal{C}$, $\hat{B}$ as the random variable associated to the bias, and $B$ the random variable associated to the extractable information regarding the bias, from $\mathcal{E}$. We can easily write the joint probability between $\hat{Y}$ and $Y$:
$p\left(Y,\hat{Y}\right)=\frac{1}{N_{C}}\left[\delta_{y\hat{y}}(1-\varepsilon_{y\hat{y}})+\bar{\delta}_{y\hat{y}}\frac{\varepsilon_{y\hat{y}}}{N_{C}-1}\right]$ (1)
@@
$p\left(\hat{Y},\hat{B}\right)=\frac{1}{N_{C}}\left[\delta_{\hat{y}\hat{b}}\rho+\bar{\delta}_{\hat{y}\hat{b}}\frac{1-\rho}{N_{B}-1}\right]$ (2)
@@
$\mathcal{I}(\hat{B},\hat{Y})=\frac{N_{B}}{N_{C}}\left[\log(N_{B})+\rho\log(\rho)+(1-\rho)\log\left(\frac{1-\rho}{N_{B}-1}\right)\right].$ (3)
The full derivation can be found in the Supp. Mat. Evidently, there is an implicit link between bias and target classes. Considering that we aim at learning a mapping between input and output, we can introduce the implicit parameter $\phi$ that encodes the tendency of our model to rely on biased features: we call this parameter model’s biasedness and it contributes as well to the error $\varepsilon$. For tractability, we will now treat the problem assuming one bias source per target class (the same is extendible to having more bias sources per target class, as well as having one same bias per multiple target classes), and for simplicity, we have ${}{N_{C}=N_{B}=N}$. Besides, we will assume that all the sources of error are due to the presence of the bias and that they are uniform across classes. It is possible to write a joint probability distribution between $\hat{B},\hat{Y},Y$, and by marginalizing over $\hat{Y}$, we can obtain
$\mathcal{I}(\hat{B},Y)=\frac{\rho+\phi(1-\rho)}{\log(N)}\log\left[\rho+\phi(1-\rho)\right]+\frac{(1-\phi)(1-\rho)}{\log(N)}\log\left[\frac{(1-\phi)(1-\rho)}{N-1}\right]+1.$ (4)
The full derivation can be found in the Supp. Mat. During debiasing, what we would like to achieve, is to minimize such a quantity, which is achieved for $\phi\rightarrow 0$: on the contrary, we can evidently see that the model is completely biased when $\phi\rightarrow 1$.
@@
$+f\left\{\frac{1}{N}\left[\frac{\phi(1-\rho)}{N-1}K_{\text{bia}}+\left(\frac{\rho^{2}(N-2)+\rho(1-\rho)(N-2)}{(N-2+\rho)}\right)\varepsilon_{\text{bia}}\right],N(N-1),N^{2}\right\},$ (6)
@@
The full derivation for (D.4) can be found in the Supp. Mat. Fig. 2c displays how the performance of a predictor drops as a function of both $K_{\text{bia}}$ and $\phi$. As we can observe, there is a clear dependence between the remotion of the bias features ($\phi\rightarrow 0$) and the final model’s performance, depending on the dataset-related parameter $K_{\text{bia}}$; if $K_{\text{bia}}$ is high, then the performance drops as $\phi$ approaches zero; on the contrary, if $K_{\text{bia}}$ is low, the performance increases as $\phi\rightarrow 0$. This tells us that debiasing, intended as fair use of features without biases, does not imply necessarily removal of these, but rather finding a proper balancing between the employed features. To achieve this, one possibility is, for example, to mask the information related to the bias, at the output of $\mathcal{E}$. Intuitively, if the information related to the bias can be disentangled from the target, this can be effectively achieved - however, if the target task intrinsically requires the employment of the information related to the bias, then its effective removal will harm the target task performance. In the next section, we will try to unveil what is the effect of removing the information related to the bias in a structured way.
@@ 3.2 Towards bias information removal
$\mathcal{I}(\hat{B},B)\leq\log\frac{\exp(s^{\mathbin{\|}})}{\frac{1}{N_{B}-1}\sum_{j}\exp(s_{j}^{\perp})},$ (7)
where $s(\cdot)$ is a similarity measure, typically expressed as cosine similarity, in the feature space (in our case, the output of $\mathcal{E}$), between samples having same bias aligned ($s^{\mathbin{\|}}$), and bias misaligned samples ($s^{\perp}$). What we aim to achieve through debiasing approaches is in general to minimize (7): we can achieve it by either minimizing the numerator (maintaining the denominator constant) or maximizing the quantity at the denominator (maintaining the quantity at the numerator constant). Some works in the literature attempted to tackle this problem directly, designing proper regularization or loss functions, requiring training of the model [55, 29, 8].
@@ 3.3 Finding the Fantastic Weights
In this section, we will provide the basics regarding how to find a mask $\mathcal{M}$ on a vanilla-trained model that removes the information related to biases. Overall, we will assume that a vanilla-trained DL model, composed of an encoder $\mathcal{E(\cdot)}$ and a task classifier head $\mathcal{C(\cdot)}$, is provided to us. We will not require any constraint regarding the way such a model is trained, and the parameters belonging to those two blocks will not be modified all along the process (apart from allowing their pruning).
Estimating bias leaking information. To estimate the amount of bias information leaking to the classifier $\mathcal{C(\cdot)}$, we will attach at the output of the encoder (the bottleneck layer) an auxiliary privacy head $\mathcal{P(\cdot)}$. We will design $\mathcal{P}(\cdot)$ such that it has the same number of parameters as $\mathcal{C}(\cdot)$: this head will be trained to estimate the upper bound on the possible information that can be used by the classifier about the bias. To this end, we can train $\mathcal{P}(\cdot)$ using a classification loss, like the categorical cross-entropy loss (CCE). Along the process, we will not allow the error signal to backpropagate through the encoder $\mathcal{E}(\cdot)$.
Metrics to remove the bias. We want to minimize the achievable performance of $\mathcal{P}(\cdot)$. Some works suggest that techniques like gradient inversion [36] are potentially employable in such a context; however, in this case, such an approach is not a good fit. More specifically, maximizing CCE does not really remove information, but simply minimizes the activation of the correct class (as also explained in some works like [43]). What we can on the contrary employ, is to minimize the mutual information between bias labels, making the output converge as much as possible to a uniform distribution. Such a term will not contribute to the update of $\mathcal{P}(\cdot)$ but it will reach the parameters in $\mathcal{E}(\cdot)$, to be then used to estimate the pruning masks.
Estimating the bias mask. In FFW we propose two possible approaches to find the pruning masks: one acts unstructuredly, while in the other we will behave in a structured manner. In an unstructured sub-network selection setup, every parameter in $\mathcal{E}(\cdot)$ will have a gating parameter $m_{i}\in\mathbb{R}$ associated with it, initialized to zero. The value of each parameter $w_{i}$ will be replaced by
@@
where $\tau$ is a temperature, $\sigma(\cdot)$ is the sigmoid function, and $\Theta(x)$ is the one-step function. To maintain the differentiability of our formulation, at the differentiation stage, we employ a straight-through estimator [9] for $\Theta(\cdot)$. As $\tau\rightarrow 0^{+}$, the value of (8) will become either $w_{i}$ (unpruned) or 00 (pruned). In the structured variant, the whole unit will be pruned, and the $j$-th neuron’s output $z_{j}$ is masked, according to
@@
All the gating parameters are initialized to zero, such that we guarantee that, at initialization, the performance of the pruned model exactly matches the one of the vanilla. Then, the proxy parameters are updated according to two different loss terms:
  • a task loss, to be kept as low as possible to maintain performance on the target task;
@@
where $\mathcal{I}\left(b,\hat{b}\right)$ is the empirical mutual information calculated for the given minibatch. The computation of the update for $m_{i}$ follows standard back-propagation rules.

### v2

@@ Abstract
Nowadays an ever-growing concerning phenomenon, the emergence of algorithmic biases that can lead to unfair models, emerges. Several debiasing approaches have been proposed in the realm of deep learning, employing more or less sophisticated approaches to discourage these models from massively employing these biases. However, a question emerges: is this extra complexity really necessary? Is a vanilla-trained model already embodying some “unbiased sub-networks” that can be used in isolation and propose a solution without relying on the algorithmic biases? In this work  [^1] Published in the European Conference on Computer Vision 2024, we show that such a sub-network typically exists, and can be extracted from a vanilla-trained model without requiring additional fine-tuning of the pruned network. We further validate that such specific architecture is incapable of learning a specific bias, suggesting that there are possible architectural countermeasures to the problem of biases in deep neural networks.
Keywords:
Debiasing Pruning Freezed model Deep learning
@@ 1 Introduction
Multiple solutions are proposed in the last luster, all relying on relatively heavy re-training approaches. More specifically, in the realm of DL debiasing, we can identify three main lines of research: (i) supervised, where the labels of the bias are provided; (ii) bias-tailored, where proxy models can capture specific biases; (iii) unsupervised, where biases are guessed directly within the vanilla model. All three approaches typically require heavy training procedures for either properly tuning the hyper-parameters of the models, or training multiple times the neural networks. This comes at another very relevant environmental cost, and besides it is not granted that such solutions can always apply to any architecture/task. Some recent studies hinder the possibility of reusing pre-trained models and removing bias sources with some adjustments [60, 61], but none of them provides a final solution yet.
In this work we show the existence of unbiased sub-networks in vanilla-trained models, providing also some analysis on the final performance. In a nutshell, we learn how to mask trained weights such that the information about the bias becomes non-extractable by the layer(s) entitled to solve the target task. Therefore, we Find the Fantastic Weights (FFW) constituting an unbiased sub-network that solves the tasks, without further fine-tuning (Fig. 1). Our results, besides providing a clear indication that such a sub-network exists in vanilla-trained overfitting models, contribute to the DL debiasing community, showing that these solutions are achievable without relying on over-complex training schemes.
@@ 2 Related works
Spurious correlations. The risk of the insurgence of spurious correlations employed for a given target task is a broadly acknowledged issue in the DL community. This phenomenon is known as shortcut learning [20], where the algorithm relies on the simplest correlations to fit the training set, but do not generalize well. Some of these can range from the texture of images [21] to biases inherent in language [23], or even sensitive variables like ethnicity or gender [44, 18]. Such behavior raises practical concerns, particularly in applications where the reliability of DL is critical, such as healthcare, finance, and legal services [13].
@@
Besides these, a new class of unsupervised approaches is rising [40], where little or even no clue on the nature of the bias is provided. Specifically, we can recognize models pre-capturing texture biases [55, 22], leveraging groups imbalances [37, 14, 3] and assuming that the bias is learned in early training stages [42, 35].
Studying impacts of neural architectures. Recently a lot of attention has been devoted in studying the influence of a deep neural network’s architecture on its generalizability. For example, Diffenderfer et al. [16] built a pruning algorithm inspired by the lottery ticket hypothesis [19] to craft compact and resilient network architectures. Similarly, Bai et al. [5] tackles a similar issue through the lenses of out-of-distribution performance. This class of approaches, although effective at the cost of enhanced computational complexity [51], does not entirely eradicate connections to spurious attributes. On the other hand, Zhang et al. and Zhao et al. suggest the efficacy of pruning weights associated with such attributes sided with some fine-tuning [59, 30], finding some counter-arguments in the literature [8] and even with works suggesting the need to employ corrective distillation terms [45]. The discussion of these empirical findings is still open.
In a recent work, it has been empirically showcased the presence of sub-networks within neural networks that exhibit reduced susceptibility to spurious features [59]. By capitalizing on the modular nature inherent in neural architectures [15], Zhang et al. were able to demonstrate that, in principle, it is possible to select, already at initialization, networks that will avoid the employment of biased features. However, this is not a sufficient condition to still guarantee the presence of such sub-networks in an already-trained model. In the era of foundation models, indeed, training a model from scratch can be very costly and impractical [58]. In this work we focus on this problem, characterizing the condition of extractability for such debiased sub-networks, drawing as well a link between bias removal and impact on the performance of the target task. We will sketch the theory motivating our approach, and we will study the problem of extracting these sub-networks, trained with vanilla strategies, without requiring finetuning.
@@ 3.1 Removing the bias impacts the performance
In this section, we will define the problem of learning a mapping between the input $\mathbf{x}$ and its target output $\hat{y}$ on a given dataset $\mathcal{D}$ (that is traditionally slit in train, validation, and test), under the assumption that there is an underlying bias $\hat{b}$ (namely, “spurious” correlations) associated to it. We can model this problem working with the random variables associated with these quantities: we define $\hat{Y}$ as the random variable associated with the target output, $Y$ as the random variable associated with the output of the DL model composed of an encoder $\mathcal{E}$ and a classifier $\mathcal{C}$, $\hat{B}$ as the random variable associated to the bias, and $B$ the random variable associated to the extractable information regarding the bias, from $\mathcal{E}$. We can easily write the joint probability between $\hat{Y}$ and $Y$:
$p\left(Y,\hat{Y}\right)=\frac{\scriptstyle 1}{\scriptstyle N_{C}}\left[\delta_{y\hat{y}}(1-\varepsilon_{y\hat{y}})+\bar{\delta}_{y\hat{y}}\frac{\scriptstyle\varepsilon_{y\hat{y}}}{\scriptstyle N_{C}-1}\right]$ (1)
@@
$p\left(\hat{Y},\hat{B}\right)=\frac{\scriptstyle 1}{\scriptstyle N_{C}}\left[\delta_{\hat{y}\hat{b}}\rho+\bar{\delta}_{\hat{y}\hat{b}}\frac{\scriptstyle 1-\rho}{\scriptstyle N_{B}-1}\right]$ (2)
@@
$\mathcal{I}(\hat{B},\hat{Y})=\frac{\scriptstyle N_{B}}{\scriptstyle N_{C}}\left[\log(N_{B})+\rho\log(\rho)+(1-\rho)\log\left(\frac{\scriptstyle 1-\rho}{\scriptstyle N_{B}-1}\right)\right].$ (3)
The full derivation can be found in the Appendix. Evidently, there is an implicit link between bias and target classes. Considering that we aim at learning a mapping between input and output, we can introduce the implicit parameter $\phi$ that encodes the tendency of our model to rely on biased features: we call this parameter model’s biasedness and it contributes as well to the error $\varepsilon$. For tractability, we will now treat the problem assuming one bias source per target class (the same is extendible to having more bias sources per target class, as well as having one same bias per multiple target classes), and for simplicity, we have ${}{N_{C}=N_{B}=N}$. Besides, we will assume that all the sources of error are due to the presence of the bias and that they are uniform across classes. It is possible to write a joint probability distribution between $\hat{B},\hat{Y},Y$, and by marginalizing over $\hat{Y}$, we can obtain
$\mathcal{I}(\hat{B},Y)=\frac{\scriptstyle\rho+(1-\phi)(1-\rho)}{\scriptstyle\log(N)}\log\left[\rho+\phi(1-\rho)\right]+\frac{\scriptstyle\phi(1-\rho)}{\scriptstyle\log(N)}\log\left[\frac{\scriptstyle(1-\phi)(1-\rho)}{\scriptstyle N-1}\right]+1.$
The full derivation can be found in the Appendix. During debiasing, what we would like to achieve, is to minimize such a quantity, which is achieved for $\phi\rightarrow 0$: on the contrary, we can evidently see that the model is completely biased when $\phi\rightarrow 1$.
@@
$\mathcal{I}$ $(\hat{Y},Y)=f\left\{{\displaystyle\frac{\scriptstyle 1}{\scriptstyle N}\left[\rho(1-\varepsilon_{\text{bia}})+(1-\phi)(1-\rho)(1-K_{\text{bia}})\right],N,N^{2}}\right\}+$
$f\left\{{\displaystyle\frac{\scriptstyle 1}{\scriptstyle N}\left[\frac{\scriptstyle\phi(1-\rho)}{\scriptstyle N-1}K_{\text{bia}}+\left(\frac{\scriptstyle\rho^{2}(N-2)+\rho(1-\rho)(N-2)}{\scriptstyle(N-2+\rho)}\right)\varepsilon_{\text{bia}}\right],N(N-1),N^{2}}\right\},$ (6)
@@
The full derivation for (0.B.4) can be found in the Appendix. Fig. 2c displays how the performance of a predictor drops as a function of both $K_{\text{bia}}$ and $\phi$. As we can observe, there is a clear dependence between the remotion of the bias features ($\phi\rightarrow 0$) and the final model’s performance, depending on the dataset-related parameter $K_{\text{bia}}$; if $K_{\text{bia}}$ is high, then the performance drops as $\phi$ approaches zero; on the contrary, if $K_{\text{bia}}$ is low, the performance increases as $\phi\rightarrow 0$. This tells us that debiasing, intended as fair use of features without biases, does not imply necessarily removal of these, but rather finding a proper balancing between the employed features. To achieve this, one possibility is, for example, to mask the information related to the bias, at the output of $\mathcal{E}$. Intuitively, if the information related to the bias can be disentangled from the target, this can be effectively achieved - however, if the target task intrinsically requires the employment of the bias information, then its effective removal will harm the target task performance. In the next section, we will try to unveil what is the effect of removing the information related to the bias in a structured way.
@@ 3.2 Towards bias information removal
$\mathcal{I}(\hat{B},B)\leq\log\frac{\scriptstyle\exp(s^{\mathbin{\|}})}{\frac{1}{N_{B}-1}{\scriptstyle\sum_{j}\exp(s_{j}^{\perp})}},$ (7)
where $s$ is a similarity measure, typically expressed as cosine similarity, in the feature space (in our case, the output of $\mathcal{E}$), between samples having same bias aligned ($s^{\mathbin{\|}}$), and bias misaligned samples ($s^{\perp}$). What we aim to achieve through debiasing approaches is in general to minimize (7): we can achieve it by either minimizing the numerator (maintaining the denominator constant) or maximizing the denominator (maintaining the numerator constant). Some works in the literature attempted to tackle this problem directly, designing proper regularization or loss functions, requiring training of the model [52, 26, 6].
@@ 3.3 Finding the Fantastic Weights
In this section, we will provide the basics regarding how to find a mask $\mathcal{M}$ on atrained DL model that removes the information related to biases. Overall, we will assume that a vanilla-trained model, composed of an encoder $\mathcal{E}$ and a task classifier head $\mathcal{C}$, is provided. We will not require any constraint regarding the way such a model is trained, and the parameters belonging to those two blocks will not be modified during the process (apart from allowing their pruning).
Estimating bias leaking information. To estimate the amount of bias information leaking to the classifier $\mathcal{C}$, we will attach at the output of the encoder (the bottleneck layer) an auxiliary bias extraction head $\mathcal{P}$ coined Privacy Head in [50] its purpose was “[to estimate] the information we desire to maintain private”: the bias . We will design $\mathcal{P}$ such that it has the same number of parameters as $\mathcal{C}$: this head will be trained to estimate the possible information that can be used by the classifier about the bias. To this end, we can train $\mathcal{P}$ using a classification loss, like the categorical cross-entropy loss (CCE). Along the process, we will not allow the error signal to backpropagate through the encoder $\mathcal{E}$.
Metrics to remove the bias. We want to minimize the achievable performance of $\mathcal{P}$. Some works suggest that techniques like gradient inversion [33] are potentially employable in such a context; however, in this case, such an approach is not a good fit. More specifically, maximizing CCE does not really remove information, but simply minimizes the activation of the correct class (as also explained in some works like [40]). What we can on the contrary employ, is to minimize the mutual information between bias labels and the output of $\mathcal{P}$, making the output converge as much as possible to a uniform distribution. Such a term will not contribute to the update of $\mathcal{P}$ but it will reach the parameters in $\mathcal{E}$, to be then used to estimate the pruning masks.
Estimating the bias mask. In FFW we propose two possible approaches to find the pruning masks: one acts unstructuredly, while in the other we will behave in a structured manner. In an unstructured sub-network selection setup, every parameter in $\mathcal{E}$ will have a gating parameter $m_{i}\in\mathbb{R}$ associated with it, initialized to zero. The value of each parameter $w_{i}$ will be replaced by
@@
where $\tau$ is a temperature, $\sigma$ is the sigmoid function, and $\Theta(x)$ is the one-step function. To maintain the differentiability of our formulation, at the differentiation stage, we employ a straight-through estimator [7] for $\Theta$. As $\tau\rightarrow 0^{+}$, the value of (8) will become either $w_{i}$ (unpruned) or 00 (pruned). In the structured variant, the whole unit will be pruned, and the $j$-th neuron’s output $z_{j}$ is masked, according to
@@
All the gating parameters are initialized to zero, such that we guarantee that, at initialization, the performance of the pruned model exactly matches the vanilla one. Then, the proxy parameters are updated according to two loss terms:
  • a task loss, wanted as low as possible to maintain performance on the target task;
@@
where $\mathcal{I}\left(b,\hat{b}\right)$ is the empirical mutual information calculated for the given minibatch and $\gamma$ is a regularization coefficient. The update for $m_{i}$ follows standard back-propagation rules.